Determination of pH of acid solution like 0.1 M HCL,0.05 M HCL,0.025 M HCL and 0.0125 M HCL by using quinhydrone and calomel electrode.compare the value of pH with the theoretically calculated values.
1
Expert's answer
2017-01-11T12:38:13-0500
The potential of quinhydrone electrode at 25 C is 0.6992 V whereas that of calomel electrode is 0.2438 V. The total potential can be written as E = E(quin. or cal.) - E(2H+/H2). Potential of the hydrogen electrode in a respective solution will be E = -0.059*lg(H+). Thus, E(quin.) = 0.6992 V - (-0.059*lg(H+)) or E(cal.) = 0.2438 V - (-0.059*lg(H+)). However, to compare these calculations with some practical results we need to have some practical results. Thus, please give your data and I will compare them.